OpenStudy (deoxna):

In the function ln(x)/sqrt(x) what does the root of the derivative mean? It's not an extrema, so what is it?

OpenStudy (anonymous):

If its not an extrema, than it is a saddle point.

OpenStudy (deoxna):

Actually, upon consulting wolframalpha, it turns out it is a maximum: https://www.wolframalpha.com/input/?i=+ln%28x%29%2Fsqrt%28x%29 But why? That would mean that the function would begin decreasing after that and would eventually have a second root, but t apparently only has one root....

OpenStudy (cruffo):

One way is to check the sign of the slope on either side of the critical value (otherwise known as cv or root). Positive slope on the left and negative slope on the right means that the cv is a maximum. Let \(f(x) = \frac{\ln(x)}{\sqrt{x}}\), then \(f'(x) = \frac{2-\ln(x)}{2x^{3/2}}\). Since \(f(x)\) is not defined at \(x=0\), we don't need to worry about the denominator of \(f'(x)\). Setting the numerator of \(f'(x)\) equal to zero, we get a solution of \(e^2 \approx 7\). To check the slope of \(f(x)\) we evaluate the first derivative: \(f'(5)\) is positive, meaning that \(f(x)\) is increasing on the left of the cv. \(f'(8)\) is negative, meaning that \(f(x)\) is decreasing on the right of the cv. So that means that the cv is a maximum :)

OpenStudy (cruffo):

Since there are no other critical values, this is the only extrema for \(f(x)\).

OpenStudy (deoxna):

Yes, but that isn’t the whole story because f that were the only change n slope, then the function would eventually cross the x-axis and hence have a second root that is not accounted for. However, the function has an inflection point at e^(8/3) so that might explain why a second root doesn't exist.